Solve by using the elimination method:
5x + 6y=5
10x + 12y=10

Answer by Theo(13342)

multiply the first equation by 2 to get:

10x + 12y = 10

your 2 equations are now:

10x + 12y = 10
10x + 12y = 10

subtract second equation from first to get:

0 + 0 = 0

since the equation is true, and all variables have canceled out, the 2 equations are identities meaning that they are really the same equation.

if you graphed both these equations, they would show up as being the same line.

take 5x + 6y = 5 and solve for y to get y = -(5/6)x + (5/6)

take 10x + 12y = 10 and solve for y to get y = (10/12)x + (10/12)

simplify the second equation to get y = (5/6)x + (5/6)

same equation will graph as the same line.

the equations are identical.

just to s how you, i will graph the 2 equations of y = (5/6)x + (5/6) and y = (10/12)x + (10/12)

see below:



both equations are identical so the graph of both of these equations looks like the same line.